Steam enters a nozzle at a pressure of 7 bar and 205°C (h = 2850 kJ /kg) and leaves at a pressure of 1.5 bar. The initial and exit velocities of steam are 40 m/s and 700 m/sec respectively. The mass flow rate through the nozzle is 1400 kg/h. the heat loss from the nozzle 11750 kJ/h. determine the final enthalpy of steam and the nozzle exit area if the specific volume is 1.24 m3 / kg.
1
Expert's answer
2020-10-25T18:28:34-0400

Enthalpy:

"H=U+PV"

"Q-W=H_1-H_2+\frac{(V_1^2-V_2^2)}{2}"

"W=P_2V_2-P_1V_1"

"H_2=H_1+P_2V_2-P_1V_1-Q+\frac{(V_1^2-V_2^2)}{2}"

"H_1=\frac{3}{2}\nu RT+P_1V_1"

"H_2=\frac{3}{2}\nu RT+P_1V_1+P_2V_2-P_1V_1-Q+\frac{(V_1^2-V_2^2)}{2}"